OpenStudy (anonymous):

solve for the variable m + 9 = -12

OpenStudy (anonymous):

m=-21

OpenStudy (fozia):

-21

OpenStudy (uri):

m =-12 -9 m = -21 ans

OpenStudy (anonymous):

-21

OpenStudy (anonymous):

m + 9 = -12 -9 -9 m=-21